OpenStudy (anonymous):

simplify the expression 4 radical 6 times radical 6

OpenStudy (aum):

\[\large \sqrt[4]{6} * \sqrt{6}~~~ \text{ or }~~~ 4 * \sqrt{6} * \sqrt{6}~~? \]

OpenStudy (aum):

\(\large \sqrt[4]{6}\) is FOURTH root of 6. \(\large 4 * \sqrt{6}\) is four times SQUARE root of 6.

OpenStudy (anonymous):

the second option except delete the multiply sign between the 4 and 6

OpenStudy (anonymous):

4 square root 6 * square root 6

OpenStudy (aum):

\[ 4\sqrt{6} * \sqrt{6} = 4 * 6 = 24 \\ \]

OpenStudy (aum):

square root x times square root x = x

OpenStudy (anonymous):

thanks how about 5 square root 24 * 3 square root